OpenStudy (anonymous):

okay so this question is confusing.. so bare with me. i am going to try to draw it out since the symbol isnt possible to make online

OpenStudy (anonymous):

\[\sqrt[3]{-3}\sqrt[3]{-5}=\sqrt[3]{15}\] true or false? explain also please

OpenStudy (anonymous):

\[\sqrt[3]{-3}\sqrt[3]{-5}=\sqrt[3]{15}\] true or false? explain also please